Solve. 9x – 0.34 = 8(x – 0.12)
I got –0.28
Is it right?

Answer by stanbon(75887)

Solve. 9x – 0.34 = 8(x – 0.12)
9x-Solve. 9x – 0.34 = 8x - 0.96
x= - 0.96+0.34
x= -0.62
================
Cheers,
Stan H.